htta//newconnectn Help Save ติ Exit ded 2 Chuck a siigie taxpayer, earns $43.500 in taxable income and $12.250 in interest from an investment in City of Heflin bonds. (Use the Required ล.Ouck eams en adotonal S24500 of taxable income, what s his marg ng lax late on the income? b. What is hi merginal rase it, instead, he had $24,500 of additional deductions
Answer #1

a

Tax on existing $43,500 taxable income is = 1,905 + 12% * (43,500-19,050) = 4,839

Tax on income after adding 24,500 additional income:

Taxable income = 43,500 + 24,500 = 68,000

Tax on $68,000 taxable income = 1,905 + 12* (68,000 - 19,050) = 7,779

Marginal tax paid on 24,500 additional income= tax on 68,000 - tax on 43,500 = 7,779 - 4,839 = 2,940

Marginal tax rate applicable = additional tax paid / additional income X 100 = 2,940/ 24,500 X 100 = 12%

Answer is 12%

b

Tax on existing $43,500 taxable income is = 1,905 + 12% * (43,500-19,050) = 4,839

Tax on income after reducing taxable income by 24,500 :

Taxable income = 43,500 - 24,500 = 19,000

Tax on $19,000 taxable income = 19,000 X 10% = 1,900

Marginal tax reduced because of additional deduction = tax on 43,500 - tax on 19,000 = 4,839 - 1,900= 2,939

Marginal tax rate applicable = additional tax deduction / additional deduction X 100 = 2,939/ 19,000 X 100 = 15.47%

Answer is 15.47%